OpenStudy (anonymous):

express sqrt -72 in i notation.

OpenStudy (anonymous):

\[\sqrt{-72} = i \sqrt{72} = \sqrt{72i} \] I think that is how you do it, not sure completely :)

OpenStudy (anonymous):

@Nurali

OpenStudy (anonymous):

you would have to know that \[\sqrt{-1} = i\] therefore